OpenStudy (anonymous):

what is the molarity of a sucrose solution (C12H22O11) if 110g of a 92.0% pure solid is dissolved per 250.0 mL of water

OpenStudy (cwrw238):

1 mole of sucrose = 144 + 22 + 16*11 = 144 + 22 + 176 = 342 g 110 g of 92% sucrose weighs = 101.2 g 101.2 in 250 ml equates to 404.8 g / litre so molarity = 404.8 / 342 = 1.184 Molar solution